How do I tell if an equation is in standard form or not? I know the standard form Ax+By=C but can you help explain this to me. Here are a couple of examples in my homework.
7y+2x=18
x+y=-8
-x+3y=-8
You do not have to answer these problems, just help me with the understanding so that I may be able to do them. I have about 15 of these and all I have to answer is if they are in standard form or not.
Thank you for any help that you may have

Answer by stanbon(75887)

How do I tell if an equation is in standard form or not? I know the standard form Ax+By=C but can you help explain this to me. Here are a couple of examples in my homework.
7y+2x=18
x+y=-8
-x+3y=-8
You do not have to answer these problems, just help me with the understanding so that I may be able to do them. I have about 15 of these and all I have to answer is if they are in standard form or not.
---------
The ones you listed are all in standard form.
They all follow the form Ax+By=C
Cheers,
Stan H.